# Task and Motion Planning in Hierarchical 3D Scene Graphs

### v1

@@ III-B Execution-Consistency
In our example, we want to prevent the agent from entering places that it should not, and so would like to show that the Inspection domain is execution-consistent with respect to the verifier $V_{place}$. Recall that $V_{place}$ can only introduce new VisitedPlace facts. As VisitedPlace does not appear in any action preconditions, the only way for a VisitedPlace fact to render a valid state plan invalid is to conflict with the goal specification. Consider the set of places that must be avoided to satisfy some goal state: $\mathcal{P}_{avoid}=\{p~{}|~{}\texttt{(not (VisitedPlace p))}\in\mathcal{G}\}$. If a place in $\mathcal{P}_{avoid}$ can only be visited by an action that explicitly lists it in the action effects, then the domain will be execution-consistent with respect to $V_{place}$. This can easily be guaranteed by preventing the motion planner from generating plans that enter places $\mathcal{P}_{avoid}\setminus\mathcal{P}_{param}$, where $\mathcal{P}_{param}$ is the set of places that appear as parameters to the action. In other words, we define a motion planner that cannot enter any Place which we might want to avoid, unless that Place is given as the parameter to the moveRelaxed action. As a result, we can be sure to never visit a potentially illegal place without (VisitedPlace p) being an effect of the action.
This notion of avoiding potentially illegal places implies something important with respect to the motion planning problem. If we identify a place to be redundant, and remove it from the planning domain, the definition of redundancy requires that any plan that would have passed through this place if we hadn’t removed it can be replicated. As a result, while the discrete PDDLStream planner does not need to be aware of pruned places, the motion planner we implement to solve for trajectories must still be able to plan paths through these places. Our motion planner must be aware of the full scene graph, regardless of what is pruned from the PDDLStream problem instance. We address this further in Sec. III-C. With that in mind, we can now consider which specific places we can remove from the PDDLStream problem instance.
Proposition 2 (Redundant Places).
Consider a problem instance in the Inspection Domain with no quantifiers that can be parameterized by a place in the goal. A place $p$ is redundant if no facts parameterized by $p$ appear in the initial or goal states, or if (not (VisitedPlace $p$)) appears as a clause in the conjunctive normal form (CNF) of the goal specification.
Proof.
First note that no actions in this domain have universal quantifiers, so we only need to check Definition 1.1 to show that a symbol is redundant. Consider a place $p$ such that (not (VisitedPlace $p$)) appears in the CNF of the goal. If $p$ parameterizes moveRelaxed, then (VisitedPlace $p$) is in the effects, violating the goal. Since moveRelaxed is the only action that can be parameterized by a place, no plan can parameterize $p$ and Definition 1.1 is trivially satisfied.
Next, consider a place $p$ that does not parameterize any initial or goal facts. For any plan $\pi$ with an action parameterized by $p$, let $a_{k}^{\mathcal{P}}$ denote an action parameterized by places $\mathcal{P}$, including $p$. Plan $\pi^{\prime}$ where $a_{k}^{\mathcal{P}}$ is replaced by $a_{k}^{\mathcal{P}\setminus p}$ is also valid, since state plans $\mathcal{I}_{\pi}$ and $\mathcal{I}_{\pi^{\prime}}$ only differ by a (VisitedPlace $p$) fact, and no action preconditions or goals involve this fact. Since $p$ was not in $\mathcal{P}_{avoid}$, the command sub-sequence corresponding to $a_{k}^{\mathcal{P}}$ is also valid for $a_{k}^{\mathcal{P}\setminus p}$. Thus, for any $\pi$ parameterized by $p$, we can construct $\pi^{\prime}$ that has an equivalent motion sequence but does not parameterize $p$, showing $p$ is redundant. ∎
We have now identified a potentially large (depending on the sparsity and geometry of the given scene graph) set of Place symbols that are redundant to our instantiation of the Inspection domain. Removing these places from our task planner’s domain, assuming the motion planner is still aware of them, should enable our solver to find valid plans that are guaranteed to have also been valid in the un-pruned problem. Moreover, the ability to prune these elements does not restrict the type of goals we are able to specify to our agent, preserving expressivity, while enabling planning at a larger scale. Our explicit method of defining our problem’s initial state is as follows:
Remark 1 (Problem Initialization).
In light of Proposition 2, we only include the following places when instantiating a problem: 1) the initial place that the robot is in and 2) any place that appears in the goal. A place $p$ that parameterizes a negated fact (not (VisitedPlace $(\texttt{not }(\texttt{VisitedPlace }p))$ that appears as a clause in the CNF of the goal specification can also be removed. Note that if (not (VisitedPlace $(\texttt{not }(\texttt{VisitedPlace }p))$ appears in the goal, but not as a standalone clause in the CNF, we do include that place. Any problem in the Inspection domain with this instantiation is guaranteed to return only plans that are execution-consistent with respect to the $V_{place}$ verifier discussed earlier, and as such, no plans need to be verified at runtime.
In implementing the resulting domain, there is a tradeoff between branching factor and required plan depth. Letting each move action be parameterized by $N$ places results in a search problem with large branching factor and expensive task planner preprocessing when running the search. In exchange, a single action can move through several places that appear in $\mathcal{P}_{avoid}$. Alternatively, we can restrict each move action to include only two places – those at the beginning and end points of the action – and preserve the set of solution trajectories. This comes at the expense of more numerous sub-problems and longer action plans, though we find this is a worthwhile trade off in the domains we consider. We can mitigate this computational cost by using an intermediate abstract planner to accelerate motion planning, which we describe below.
III-C Motion Planning Through Places
Critical to the success of our approach is the implementation of a motion planner that is capable of planning over large distances, while avoiding specific regions as specified by the task planner. Solving for a valid plan for a TAMP problem in a large environment requires solving for trajectories between potentially many different possible start and end poses, so the efficiency of the motion planner has a large effect on the overall runtime of the planner. The connectivity information between places in the scene graph is a useful tool for focusing the search space for feasible trajectories over large distances, since each place represents a region of navigable space. However, the connectivity information in the scene graph does not account for robot kinematic constraints or necessary standoff distances from obstacles. Thus, though we leverage information from the scene graph when possible, we still need to run a kinematic motion planner.
To find a motion plan between two configurations $c_{1}$ and $c_{2}$, we plan through the places layer of the complete, unpruned scene-graph, using A* to find a sequence of places $[p_{1},...p_{N}]$ with corresponding centers $[x_{1},...x_{N}]$ such that $c_{1}$ is in place $p_{1}$, $c_{2}$ is in $p_{N}$, and edge $(p_{i},p_{i+i})$ is present in the scene graph. Though we have pruned the PDDL planning domain to accelerate the discrete component of search, we still retain the full information in the scene graph in order to generate trajectories. This is a much simpler search problem than using the PDDL domain of the full scene graph. Second, a kinematic planner plans a path through these positions. This path ignores obstacles that may be along the path and does not necessarily stay within places on the reference path, so it is very fast to generate. For a robot with a kinematic bicycle model, this initial path might be a series of Dubins curves calculated between sequential place centers. Next, this path is checked for collisions against the geometry of all objects in the scene. Any segments of the trajectory that are rendered infeasible by collision with objects or overlap non-traversable places are re-solved by a planner that considers the obstacles, such as RRT [14]. This two-phase approach allows for fast solving of “easy” portions of the trajectory, while being able to spend time searching for feasible solutions to more difficult parts of the path. Better alignment between which edges are present in the scene graph and kinematic feasibility for the robot leads to better performance of this heuristic.
This coupling of the motion planner with the higher-level discrete planner using the places layer of the scene graph, gives us in actuality a three level hierarchical planner: 1) an abstract discrete task planner for interacting with the environment using the PDDL domain based on the pruned scene graph described in Sec. III-A, 2) a navigation planner for moving from place to place in the places layer of the full scene graph, and 3) a low-level motion planner using the abstract plan and the geometry of the scene graph to produce kinematically feasible trajectories. This planner architecture aligns naturally with the scene graph hierarchy, and we notice that with this factorization, each planning problem we care about becomes considerably simpler. The top-level planner only needs the subset of relevant the places layer, so we get a speedup by eliminating the parts of the scene graph that are redundant and reducing the planning horizon. The abstract navigation planner only plans through the places layer to one sub-goal at a time, and so can take advantage of Euclidean distance heuristics. Finally, the fine-grained motion planner is guided by the solution of the plan through the places layer, and so does not suffer from the usual challenges of finding motion plans over large distances. Crucially, we are able to do this without restricting the types of goals we can specify to the planner.
III-D Weakly Redundant Symbols and Object Pruning
We have demonstrated the ability to identify elements of a planning instance that are provably redundant when searching for a plan. However, there may be cases where planning could be made vastly more efficient by pruning elements of a problem that do not meet the strong condition of Definition 1. Consider, for example, the case where a robot is tasked with travelling across a large scene graph in the Inspection domain, containing many suspicious objects. The robot can inspect and neutralize an obstacle to pass it safely, but depending on the geometry of the scene, there may exist a path to the goal that is completely free of any obstacles. Alternatively some subset of the objects may be needed to be cleared to find safe passage. Regardless of that reality, none of the objects in the scene meet our definition of redundancy due to the fact that motion sequences which involve inspecting an object cannot be generated once the object is removed by the domain. The conditions for a symbol being redundant are quite strong – every trajectory must be preserved when removing the symbol. We now introduce a weaker notion of redundancy.
Definition 4 (Weakly Redundant Symbols).
A symbol x is weakly redundant if both of the following hold:
  1. 1.
    For every plan $\pi$ where x parameterizes an action, there is another plan $\pi^{\prime}$ that is valid where x is not an action parameter.
  2. 2.
    No action preconditions or goals, expressed in negative normal form, contains a universal quantifier that can be parameterized by x.
In contrast to strongly redundant symbols, removing a weakly redundant symbol from a planning instance may limit the set of motion sequences that can be found. An object is weakly redundant if a valid plan can be found without that object being necessary to the plan (e.g., if the robot can find a path around the object in question). Because the applicability of this definition may depend on the geometry of the scene, it is more difficult to identify symbols with weak redundancy, though it potentially applies to a wider set of objects. Consider two sets of symbols, $\mathcal{O}_{1}\subseteq\mathcal{O}_{2}\subseteq\mathcal{O}$. If there is a valid plan $p$ that is only parameterized by objects in $O_{1}$, then the objects in $O_{2}\setminus O_{1}$ are weakly redundant and do not need to appear in problem instance in order to find a valid plan.
@@ III-D Weakly Redundant Symbols and Object Pruning
Unfortunately, identifying weakly redundant objects in the Inspection domain can be as hard as solving the original planning problem. However, we can still take advantage of the sparsity that exists in many scene-graphs by considering an incremental approach for identifying a subset of objects without (some) weakly redundant objects. We begin planning by including some subset of all objects, and attempt to solve the planning problem. If this limited problem has a valid solution which is also a valid solution to the original problem, then we have found a plan (and have shown the excluded objects were in fact weakly redundant). Otherwise, we incrementally add objects to the planning problem, and repeat (Algorithm 1).
Identifying the initial set of objects $\mathcal{O}_{S}$ is not trivial. These objects are those that we are certain are neither redundant, nor weakly redundant. We have proven in the Inspection domain to be able to identify Places from the scene graph which are not redundant as described in Remark 1, so those are added to the problem. Any Object in the scene graph is potentially weakly redundant, depending on scene-geometry, unless it is explicitly mentioned in the goal, so we only add those goal Objects to the initial problem instance. For general problems, there is a body of literature dedicated to identifying object relevance, either by a reachability analysis [4, 5], or otherwise learning to predict importance [19] which could also be used to augment this initial set.
We then initialize a problem instance with these objects $\mathcal{O}_{I}=\mathcal{O}_{S}$. Let $\mathcal{O}_{R}$ denote the rest of the objects in the scene graph. Next, in accordance with the adaptive algorithm in Garrett et al. [6], we optimistically assume that certain sub-problems are solvable, and identify candidate action-plans which, if this assumption were true, would be a valid plans. We refer to these candidates as a plan skeletons. The set of plan skeletons for the problem instance involving objects $\mathcal{O}_{I}$ are generated by GetSkeletons (Line 7). SolveSubProblems (Line 8) attempts to solve the relevant sub-problems in a given candidate plan skeleton. If a sub-problem fails to be solved, we query the solver for feedback information (which objects in the full set $\mathcal{O}_{R}$ are a potential cause of failure). If all sub-problems were solved successfully, we return the full valid plan. Otherwise, we choose to add one or more new objects to the problem with GetNewObjects, which considers the feedback about the failure mode that CheckSolution (Line 9) returned. This approach will continue to incrementally add objects until a plan is found, and so is complete [^1] Assuming SolveSubProblems is complete and GetSkeletons will return arbitrarily long skeletons once all objects have been added..
@@
1:procedure IncrementalObjectSolver($A,S,I,G$)
2:     $\mathcal{O}_{S}\leftarrow GetRelevantObjects(\mathcal{O})$triangleright Find objects relevant for non-collision reasons
@@
6:         $SkeletonInfo=[~{}]$
7:         for $k\in GetSkeletons(A,S,\mathcal{O}_{I},G)$ do
@@
9:              $Feedback\leftarrow CheckSolution(T,\mathcal{O}_{R})$
10:              $SkeletonInfo.\texttt{append}(Feedback)$
@@
12:         if $\pi\in SkeletonInfo$ is valid then
@@
Existing literature related to identifying irrelevant objects in a planning problem often focuses on ignoring or more efficiently encoding symbols based on careful analysis of the PDDL problem’s logical factorization  [10, 5]. In contrast, our algorithm is well suited to address the geometric dependence of movement actions on objects that may render a trajectory infeasible. A movement action must avoid collision with all objects in the environment, and it is difficult to show that an object cannot interfere with any plan and can be ignored without explicitly inspecting the geometry. Algorithm 1 can be seen as a way of lazily uncovering which of these objects do interfere with plans that would otherwise be valid.
We take advantage of the hierarchy in the scene graph to implement GetNewObjects, which is tasked with identifying which new objects to add to the planning problem based off feedback from failed sub-problems. Each time we attempt to solve a motion plan, we produce a plan that avoids the objects that have been added to our planning domain (if one exists). Then, we check that plan against the full set of objects. If one of those were intersected, and we cannot find a plan that does not intersect any of the full set of objects, the first object is potentially relevant to the planning problem, and we add it to the planning domain. By adding this object to the planning problem, the planner can now consider inspecting and neutralizing it, thus introducing a new potential path to a goal state. We envision many potential extensions to this approach, such as learning parameters for these subroutines, which we do not explore in this work. In the next section we evaluate our planning approach in the Inspection domain.
IV Evaluation
@@
Figure 2: Three of the maps used for evaluation (not including the KITTI environment). Top Left: narrow alley map. Top Right: A simple 10x10 grid world map. Bottom: Scene graph built from real data collected by a robot in an office environment.
@@
There are three primary axes of complexity for planning problems in large scene graphs that we are interested in investigating. One is the complexity of the goal specification, another is correlated to the scale of the environment, and the third is related to the geometry of the scene (and potential obstructions therein). In our experiments, we aim to identify which types of planning problems our planning approach is well suited for considering these sources of complexity. To that end, we compare our encoding of the Inspection domain to the dense, direct encoding in a variety of different settings. We test on four map archetypes (Fig. 2)– a synthetic small constrained alleyway, a synthetic 10x10 gridworld, a scene graph built from real data in an office environment comprising 557 Places and 28 Objects, and a much larger scene graph built from the KITTI dataset composed of 17861 Places and 1315 Objects (Figs. 1 and 2). For each environment, we test several different goal clauses across different variations in robot and object initial conditions.
@@
To randomize tasks across trials, we define a mechanism for sampling goal specifications according to an increasing number of clauses in Disjunctive Normal Form (DNF). More explicitly, we generate goals in DNF with N clauses, where each clause has K conjunctions. For example, if N = 2 and K = 3, the goal looks like this: (Or ($C_{1}$, $C_{2}$)), where we sample 3 clauses for $C_{i}$, e.g., (And ((Visited P1), (Safe O4), (Not (Visited P9)))).
IV-A Scene Graph Size
Figure 3: This log-log plot compares the time (in seconds) to solve tasks of comparable complexity across different environmental scales for the dense formulation and the proposed sparse formulation. In the smallest environment, the dense baseline planner performs just as well as our approach. However, as we increase the scale, first to a 10x10 grid, and then to a real scene graph constructed on the floor of an academic building, we see the benifits of our approach.
@@ IV-A Scene Graph Size
First, we investigate the effect of scene graph size on our ability to plan to a set of goals of consistent complexity. For this set of trials, our goal complexity is (N, K) = (3, 3), and we compare planning time for the direct encoding to our planner, highlighted in Fig. 3. Each point on the scatter plot corresponds to a single trial and, as highlighted in the legend, different colors correspond to different environment types. We begin in the Alley environment which, as shown in Fig. 2, is a 2x6 grid world containing only connected places. In this environment, for this level of complexity, our planner performs about as well as the dense encoding, as there is not much advantage to sparsification in such a small environment. As we scale up however, this changes. The next environment we consider is a 10x10 grid, where we see modest improvement as shown in Fig. 3. In the figure, any samples above the black line indicate that our planner outperforms the dense baseline. As we scale up even further, into the indoor DSG, we see the baseline planner taking on the order of hundreds of seconds to plan on average, while our planner averages in the tens of seconds. We note that we only considered goals that involved visiting or not visiting certain Places in the graph. When we attempted to introduce object inspection, the dense planner timed out before finding a plan in all instances. Similarly, when testing the baseline in the KITTI scene graphs, it was also unable to find solutions for goals of any complexity. Planning with these relatively simple goals, our planner experienced only a modest increase in planning time as the size of the map scaled.
IV-B Goal Complexity
Figure 4: Here we highlight how our approach scales with the complexity of the goal specification in the simple 10x10 grid world. As we increase the number of unique PDDL objects in the goal specification, the problem is no longer sparse, and so no longer benefits from our approach. Eventually, the dense, direct encoding begins to outperform our method.
@@ IV-B Goal Complexity
Next, we consider the effect of increasing goal complexity on planning time. To do this, we investigate a series of various different goal constructions in the Grid World environment. Specifically, we run experiments for K = 5 and K = 10, incrementing the N value from 1 to 5 for each K value. Stated differently, we have conjunctive clauses that are either 5 or 10 atoms long, and “Or” together 1, 2, 3, 4, or 5 of these conjunctions. Each of these clauses is once again either (Visited P1) or (Not (Visited P1)). Figure 4 presents a plot comparing the complexity of the goal in terms of total unique symbol referenced vs planning time. For less complex goals in this environment, our planner outperforms the dense planner, up to a crossover point at around 20 unique objects. At this point, the benefits of sparsity disappear for a scene of this size. Notably however, if we were to introduce obstructing objects, the dense direct encoding baseline is unable to solve goals of any meaningful complexity. Once again, not included in this plot are the results on the KITTI dataset, as the direct encoding never successfully completes a trial in this setting due to timing out.
IV-C Object Obstruction in KITTI
Figure 5: An example plan from the KITTI environment. The robot begins in the bottom left corner and is tasked with inspecting one object (denoted by the red triangle at the end of the black trajectory). Along the way, there are numerous objects potentially blocking the agent’s path, so it must add at least one to its planning domain. After inspect and neutralizing this object, the robot is free to reach its goal.
@@ IV-C Object Obstruction in KITTI
Finally, we investigate the performance of the incremental object solver algorithm in task instances where objects not directly listed in the goal must be inspected in order to solve the task. In this experiment, we give a robot one of two goal types in a scene graph built from the KITTI dataset (Figs. 1 and 5): either (Visited $P_{i}$) or (Safe $O_{j}$). Given the size of the map, satisfying these goals may require the agent to traverse a large distance, but more importantly, if there are obstructing objects in the way, it may be forced to inspect and neutralize them to find a safe path to the goal. As a baseline, we sample 20 goals in the map shown in Fig. 5 using our planner without any of the objects being labeled as suspicious. In this case, they do not obstruct the agent’s path, and we find plans in 19 of 20 trials.
@@
Next, we “activate” 13 objects in the scene by labeling them as suspicious. A suspicious object has an inflated radius that is only safe for the robot to enter after it has been inspected and neutralized (in the KITTI scene, this radius is large enough to block an entire road as shown by the blue objects in Fig 5). For the agent to inspect the object, it has to sample a pose that is traversable and within line-of-sight and range of the object. Then, by taking the inspect action, the object is now safe, and can be passed. To demonstrate the difficulty of this task, and to highlight the importance of object pruning, we attempt to solve these same tasks without using our incremental feedback approach for pruning potentially weakly redundant objects (Sec. III-D). Instead, we add all suspicious objects to the scene directly. Using this encoding, the planner only succeeds in finding a plan in 4 out of 20 trials. Inspecting these solutions further reveals that in all 4 of these successful cases, there was a direct path to the goal without inspecting any objects. This result makes sense, as the odds of sampling the correct object to inspect is low without the benefit of geometric information.
@@
Finally, we test our proposed approach with pruning objects from the planning domain, and incrementally adding them back if they impacted the failure of a sub-problem during search (Sec. III-D). Our planner is able to solve 12 of the 20 trials, including 9 cases wherein the agent inspected one or more obstructing objects on the way to its goal. These experiments further demonstrate the importance of our proposed approach to sparsifying otherwise dense, long-horizon planning problems. An example plan, where the agent investigates two objects on the way to its goal is shown in Fig. 5.
V Related Work
There has substantial recent work enabling construction of information-rich scene graphs. Wu et al. [23] use a graph neural network to build higher-level abstractions from sensor data. Hughes et al. [12] and Bavle et al. [3] build scene graphs while localizing a robot’s motion, allowing for co-optimization of pose and scene estimation. Other recent directions include the registration of scene graph pairs [17].

### v2

@@ 3.3 Execution Consistency
In our example, we want to prevent the agent from entering places that it should not, and so we should show that the Inspection domain is execution consistent with respect to the verifier $V_{place}$. Recall that $V_{place}$ can only introduce new VisitedPlace facts. As VisitedPlace does not appear in any action preconditions, the only way for a VisitedPlace fact to render a valid state plan invalid is to conflict with the goal specification. Consider the set of places that must be avoided to satisfy some goal state: $\mathcal{P}_{avoid}=\{p~{}|~{}\texttt{(not (VisitedPlace p))}\in\mathcal{G}\}$. If a place in $\mathcal{P}_{avoid}$ can only be visited by an action that explicitly lists it in the action effects, then the domain will be execution consistent with respect to $V_{place}$. This can easily be guaranteed by preventing the motion planner from generating plans that enter places $\mathcal{P}_{avoid}\setminus\mathcal{P}_{param}$, where $\mathcal{P}_{param}$ is the set of places that appear as parameters to the action. Our mid- and low-level motion planners are constrained not to enter a place which we might want to avoid, unless that place is given as the parameter to the moveRelaxed action, ensuring execution consistency. Note that the verifier need not be actually implemented, but the concept can be used to prove execution consistency.
3.4 Ignoring Irrelevant Objects
We have demonstrated the ability to identify and ignore elements of a planning instance that are redundant when searching for a plan. However, many symbols are not redundant according to our definition, but might still be safely ignored. Objects which might obstruct motion, for example, are not redundant because an “inspect” motion primitive will never be generated for an object that has been removed from the planning instance. Nevertheless, there are clearly cases when an agent can ignore objects when planning, such as an object that is not part of an agent’s goal and is far from the agent’s path to the goal. As with the places, ignoring objects can accelerate planning by reducing the branching factor at the highest level of search. In contrast to places however, the objects that should be ignored cannot be identified from the logical structure of the planning instance alone; the problem geometry must also be considered.
@@ 3.4 Ignoring Irrelevant Objects
We propose an incremental approach to identifying relevant symbols. We begin by including some subset of all symbols $\mathcal{O}_{S}$ in the domain of the high-level planner, and attempt to solve the planning problem. If this limited problem has a valid solution which is also a valid solution to the original problem, then we have found a plan. Otherwise, we incrementally add symbols to the planning problem, and repeat (Algorithm 1). The inner loop (Line 7) corresponds to solving a TAMP instance with symbols $\mathcal{O}_{I}$. The outer loop (Line 5) corresponds to adding more symbols to $\mathcal{O}_{I}$ when we fail to find a solution.
The performance of this incremental planning approach depends on three key choices: which initial symbols are chosen in $\mathcal{O}_{S}$, when new symbols are added to the planning instance, and how the new symbols $\mathcal{O}_{new}$ are chosen. As long as all symbols are eventually added to the planning instance, this planning approach will maintain the completeness properties guaranteed by the chosen PDDLStream solution algorithm [9].
@@
1:procedure IncObjSolver($A,S,\mathcal{O},I,G$)
2:     $\mathcal{O}_{S}\leftarrow GetRelevantObjects(\mathcal{O})$
@@
6:         $SkelInfo=[~{}]$
7:         for $k\in Skeletons(A,S,\mathcal{O}_{I},G)$ do
@@
9:              $Feedback\leftarrow Check(T,\mathcal{O}_{R})$
10:              $SkelInfo.\texttt{append}(Feedback)$
@@
12:         if $\pi\in SkelInfo$ is valid then
@@
The initial set $\mathcal{O}_{S}$ should be as small as possible while still including the symbols necessary to find a plan. In particular, we can begin by including symbols based on the problem’s logical structure. For the Inspection domain, we include the non-redundant places identified in Remark 3.1 and any objects that appear in the goal. In general, there is a large body of literature dedicated to identifying object relevance, such as by reachability analysis [5, 8] or learning to predict importance [24], which may identify more symbols to add to the initial set.
We must decide how many task plan skeletons will be checked by $Skeletons$ and how much time will be spent attempting to solve the continuous subproblems before adding new symbols to the planning instance. In the Inspection domain, the full problem only has a solution if the pruned problem has a solution, so we choose to stop iterating through plan skeletons once we find a solution to the reduced problem. In general, a maximum time must be set for iterating through plan skeletons (we do so according to the Adaptive approach [9]). Finally, the choice of symbols to add to the planning instance can be informed by feedback ($Check$, Line 9) from failed solutions to subproblems. In our domain, we add objects which block the robot’s motion on an otherwise-feasible trajectory.
4 Evaluation
@@
Figure 3: Three of the maps used for evaluation (not shown is the KITTI or building 45 environments). A narrow alley map, a simple 10x10 grid world map, and a scene graph built from real data collected by a robot in an office environment.
@@
We characterize how our method’s performance depends on goal complexity, environment scale, and scene geometry. We compare our encoding of the Inspection domain to the dense, direct encoding in a variety of different settings. We test on four map archetypes (Fig. 3) – a synthetic small constrained alleyway, a synthetic 10x10 gridworld, a scene graph built from real data in an office environment comprising 557 Places and 28 Objects, and a much larger scene graph built from the KITTI dataset composed of 17861 Places and 1315 Objects (Figs. 1 and 3). For each environment, we report planning times for several different goal clauses across different variations in robot and object initial conditions. The planning time includes the incremental detection of relevant objects and the PDDL solver’s preprocessing of the planning problem. The time to convert a scene graph into the planning problem is insignificant compared to the planning time. Finally, we also implement and test our planner on a Spot robot.
@@
To randomize tasks across trials, we define a mechanism for sampling goal specifications according to an increasing number of clauses in Disjunctive Normal Form (DNF). A goal with complexity (N, K) is a formula in DNF with N clauses, where each clause has K conjunctions. For example, for complexity (2, 3), the goal has the form (Or ($C_{1}$, $C_{2}$)), where $C_{i}$ is a clause consisting of three facts e.g., (And ((Visited P1), (Safe O4), (Not (Visited P9)))).
@@
Scene Graph Size: First, we investigate the effect of scene graph size on our ability to plan. For this set of trials, the goal complexity is (N, K) = (3, 3), and we compare planning time for the direct encoding to the planning time for our planner, shown in Fig. 4A. Each point on the scatter plot corresponds to a single trial and different colors correspond to different environment types. Any samples above the black line indicate that our planner outperforms the dense baseline. In the small Alley environment, our planner performs about as well as the dense encoding, as there is not much advantage to sparsification in such a small environment. As we scale up however, the relative performance of our planner improves. In the 10x10 grid, we see modest improvement as shown by the red points in Fig. 4A. As we scale up even further, with the small-scale scene graph, we see the baseline planner taking hundreds of seconds to plan, while our planner averages in the tens of seconds. This experiment only considers goals that involve visiting or not visiting certain places in the scene graph. When we attempted to introduce object inspection, the dense baseline planner timed out before finding a plan in almost all instances. Similarly, when testing the baseline in the KITTI scene graphs, it was also unable to find solutions for goals of any complexity. Our proposed planner experienced only a modest increase in planning time as the size of the map scaled.
@@
Goal Complexity: Next, we consider the effect of increasing goal complexity on planning time. To do this, we investigate a series of different goal constructions in the Grid World environment. Specifically, we run experiments with goal complexity (N, K), for K = 5 and K = 10, and N from 1 to 5. Goal facts are chosen to be either visiting or not visiting specific places. Fig. 4B presents a plot comparing the complexity of the goal in terms of total unique symbol referenced vs planning time. For less complex goals in this environment, our planner outperforms the dense planner, up to a crossover point at around 20 unique objects. Advantages from the additional structure in the dense formulation outweigh the gains of our sparser method when a large percentage of the place symbols are relevant to the goal. The direct encoding never successfully completes a trial in the KITTI dataset due to timing out.
Figure 4: A) Comparison of the time to solve tasks of comparable complexity across different environmental scales for the dense formulation and the proposed sparse formulation. B) The scaling of our approach with the complexity of the goal specification in the simple 10x10 grid world. As we increase the number of unique PDDL objects in the goal specification, the problem is no longer sparse, and so it no longer benefits from our approach.
@@
Object Obstruction: Next, we investigate the performance of the incremental object solver algorithm in task instances where objects not directly listed in the goal must be inspected in order to solve the task. In this experiment, we give a robot one of two goal types in a scene graph built from the KITTI dataset (Figs. 1 and 5): either (Visited $P_{i}$) or (Safe $O_{j}$). Given the size of the map, satisfying these goals may require the agent to traverse a large distance, but more importantly, if there are obstructing objects in the way, it may be forced to inspect and neutralize them to find a safe path to the goal. As a baseline, we sample 20 goals in the map shown in Fig. 5 using our planner without any of the objects being labeled as suspicious. In this case, they do not obstruct the agent’s path, and we find plans in 19 of 20 trials.
@@
Next, we “activate” 13 objects in the scene by labeling them as suspicious. A suspicious object has an inflated radius that is only safe for the robot to enter after it has been inspected and neutralized (in the KITTI scene, this radius is large enough to block an entire road as shown by the blue objects in Fig 5). For the agent to inspect the object, it has to find a pose that is traversable and within range of the object. Then, by taking the inspect action, the object becomes safe, and can be passed. To highlight the importance of object pruning, we attempt to solve these same tasks without using our incremental feedback approach for object pruning (Sec. 3.4). Instead, we add all 13 suspicious objects to the scene directly. Using this encoding, the planner only succeeds in finding a plan in 4 out of 20 trials. Inspecting these solutions further reveals that in all 4 of these successful cases, there was a direct path to the goal without inspecting any objects. This result makes sense, as the odds of sampling the correct object to inspect is low without the benefit of geometric information.
@@
Finally, we test our proposed approach of incrementally adding objects to the planning instance (Sec. 3.4). Our planner solves 12 of the 20 trials, including 9 cases wherein the agent inspected one or more obstructing objects on the way to its goal. Failure to find plans is caused by PDDLStream not successfully finding sequences of inspection poses on the correct side of obstructing objects when several such objects needs to be inspected to find a plan. These experiments further demonstrate the importance of our proposed approach to sparsifying otherwise dense, long-horizon planning problems. An example plan, where the agent investigates two objects on the way to its goal is shown in Fig. 5.
Figure 5: An example plan from the KITTI environment. The robot begins in the top left, and is tasked with inspecting one object (denoted by the red triangle at the end of the trajectory). Along the way, there are numerous objects potentially blocking the path, so we must add at least one to its planning domain. After inspecting and neutralizing this object, the robot can reach its goal.
Real-World Manipulation: Finally, we demonstrate our planner in a real-world setting (Fig. 2), using a Boston Dynamics Spot quadruped to build a 3D scene graph in real-time in a university building. In order to demonstrate that our approach is effective on domains different from “Inspect”, we implement a “Retrieval” domain, which adds additional Pick and Place actions to the Inspection domain, enabling the robot to move objects around the environment. The VisitedPlace predicate is evaluated from the robot’s body position, so that the manipulator’s swept volume does not need to be considered when reasoning about places to avoid. The goal specifies which place an object should be in (e.g., (ObjectAtPlace O1 P3), denoting a goal state where object O1 is in place P3). For each trial we run, we structure the environment such that there is an obstruction preventing the robot from reaching its target object. To solve the task, the robot must move this obstruction out of the way, before retrieving the specified object. We encoded pick, place, inspect, and move skills for the robot.
Like the KITTI domain, we rely on our incremental object adding approach to only add the obstructing object to our high-level planner. Scattered throughout the environment are a number of different objects (Fig. 2), which would lead to an intractable problem if we were to consider them all. We ran ten trials, each time finding plans, and highlight several successful executions in the video supplement [^2] https://youtu.be/v8fkwLjBn58. While the planner reliably finds feasible plans, execution often requires several attempts due to failures when executing the Pick skill and Spot’s local planner failing in certain constrained passages.
5 Related Work
There has been substantial recent work enabling construction of information-rich 3D scene graphs, initially introduced by Armeni et al. [2]. Following works have focused on constructing 3D scene graphs from real-world sensor data [23], real-time performance [15, 4], and improving the higher-level abstractions [30, 26]. The strong performance of foundation models on open-vocabulary tasks has led to a series of works on combining open-vocabulary language embeddings with 3D scene graphs [12, 21, 29]. These open-vocabulary works all feature object navigation or retrieval tasks executed on real robots, although the task structure is simple and the focus is on mapping natural language to an object grounded in the scene graph.